Two identical sinusoidal waves with wavelengths of 1.5 m travel in the same direction at a speed of 10 m/s. If the two waves originate from the same starting point, but with time delay ∆t between them, and with resultant amplitude A_resultant = √3 A then ∆t will be equal to:


0.00625 sec


0.0125 sec


0.025 sec


0.005 sec


0.01 sec

Clear selection



1
Expert's answer
2021-05-25T14:25:44-0400

Let "\Delta t" be the required time delay or time difference

Phase Difference "\phi = \dfrac{2\pi}{T}\Delta t"

Given,

Velocity of wave v = 10 m/s

Wavelength "\lambda" = 1.5 m


Now,

"v = f\cdot \lambda\\\Rightarrow f=\dfrac{v}{\lambda}=\dfrac{10}{1.5}=\dfrac{20}{3}"


So, "\dfrac{1}{T}=\dfrac{20}{3}"


Therefore Phase Difference "\phi = \dfrac{2\pi}{T}\Delta t = 2\pi \dfrac{20}{3}\cdot \Delta t=\dfrac{40\pi}{3}\Delta t"


For Resultant Amplitude,

"A_r=\sqrt{A_1^2+A_2^2+2A_1A_2cos\phi}"

In this case "A_1=A_2=A"

So,

"A_r=\sqrt 2A\sqrt{1+cos\phi}\\A_r=2Acos\frac{\phi}{2}"


Given, Resultant amplitude "A_r=\sqrt 3 A"

So,

"\sqrt3 A =2Acos\frac{\phi}{2}\\cos\frac{\phi}{2}=\dfrac{\sqrt 3}{2}\\\dfrac{\phi}{2}=\dfrac{\pi}{6}\\\Rightarrow\phi=\dfrac{\pi}{3}"

"\implies\phi=\dfrac{\pi}{3}=\dfrac{40\pi}{3}\Delta t\\\Rightarrow\Delta t =\dfrac{1}{40}s=0.025\ \ sec"


Answer: "\boxed{\Delta t=0.025\ \ sec}"